Hi, can I have some help with this problem?
A 10-ft-long stem of bamboo is broken in such a way that its tip touches the ground t = 3 ft from the base of the stem What is the height of the break? (Hint: Use the Pythagorean Theorem.)

Answer by scott8148(6628)

the standing portion of the stem forms a right angle with the ground
__ the broken portion is the hypotenuse of a right triangle

let x=height, so 10-x=hypotenuse (broken portion)

x^2+3^2=(10-x)^2 __ x^2+9=100-20x+x^2 __ 20x=91